# SPYGLASS. VII-A. The Demographics and Ages of Small Nearby Young Associations

### v1

@@ Abstract
Recent Gaia-based young stellar association surveys have revealed dozens of low-mass populations that have, until recently, been too small or sparse to detect. These populations represent a largely unstudied demographic with unknown origins, and their relative isolation may minimize gravitational disruptions that impact traceback, making them compelling targets for dynamical studies. In this paper, we survey 15 of these isolated young associations for the first time: Andromeda South (SCYA-97), Aquila East, Aries South (SCYA-104), Cassiopeia East (SCYA-43), Canis Major North, Leo Central (SCYA-2), Leo East (SCYA-3), Theia 72, Ophiuchus Southeast, Scutum North (SCYA-70), Taurus-Orion 1 (TOR1), Theia 78, Vulpecula East (UPK 88), SCYA-54, and SCYA-79. By combining Gaia astrometry and photometry with new ground-based spectroscopic measurements, we assess the membership of each population, search for substructure, analyze their demographics, and compute ages. We find that the smallest populations in our sample contain <20<20 $M_{\odot}$ of stellar mass, making them the smallest associations ever detected. Four host substantial substructure, including TOR1, where we discover TOR1B, a new 9 $M_{\odot}$ association with radial velocities inconsistent with an origin in the parent complex. Using PARSEC isochrones, we produce self-consistent ages for all populations supported by dynamical and lithium depletion ages, which range from 6.9 $\pm$ 0.5 Myr in TOR1 to 42.8 $\pm$ 2.4 in AndS. Our results provide the first detailed overview of the properties of these populations, characterizing a largely unknown category of young associations that may have an important role in tracing the processes that guide local star formation.
@@
facilities: Gaia, McDonald Observatory 2.7m Harlan J. Smith Telescope, Sloan Digital Sky Survey BOSS and APOGEE spectrographs at APO and LCOsoftware: astropy (Astropy Collaboration et al., 2013, 2018, 2022), numpy (Harris et al., 2020), pandas (pandas development team, 2020), LMFIT (Newville et al., 2016) saphires (Tofflemire, 2019), matplotlib (Hunter, 2007)
@@ 1 Introduction
Most nearby young stars are not found in isolation, but rather in co-moving star clusters and associations (de Zeeuw et al., 1999; Krumholz et al., 2019). These populations inherit motions from their parent molecular cloud, and as such their positions, velocities and ages produce a record of the positions and dynamics of star-forming clouds long after their dispersal (e.g. Galli et al., 2023; Swiggum et al., 2024). On small scales, young stellar populations are used to detect star formation patterns, such as sequential and triggered star formation, where feedback from early-forming stars or gas collisions may compress gas and trigger new stellar generations (Tan, 2000; Inutsuka et al., 2015; Posch et al., 2023). On larger scales, the locations of these clouds at formation collectively trace the sites of dense gas in our Galaxy over time. An improved census of recent star formation will therefore reveal not just local star formation patterns, but also kpc-scale gas patterns that trace larger structures like spiral arms.
@@
The data releases of the Gaia spacecraft are making the local star formation record more complete than ever before (Gaia Collaboration et al., 2016a, 2022). Recent Gaia-based surveys have substantially deepened our record of nearby associations, ranging from large, heavily substructured stellar complexes to small, seemingly isolated associations (e.g. Kounkel & Covey, 2019; Hunt & Reffert, 2023). Many of the larger populations in the solar neighborhood have already been studied, often revealing extensive substructure and multi-generational star formation patterns (e.g., Cantat-Gaudin et al., 2019; Kerr et al., 2021; Briceño-Morales & Chanamé, 2023; Ratzenböck et al., 2023; Sánchez-Sanjuán et al., 2024; Kerr et al., 2024).
@@
However, associations containing less than a few hundred solar masses often lack dedicated studies of their structure, demographics, and ages, especially outside of the nearest 100 pc (de Zeeuw et al., 1999; Gagné & Faherty, 2018). As such, we know little about their typical dynamics and star formation patterns. We also know little about whether they emerge from the kpc-scale gas structures in the solar neighborhood or any clouds that may have preceded their assembly, or whether they form in truly isolated environments. If their formation is isolated, they may be better-suited for precise dynamical studies, having had minimal interactions with particularly massive nearby gas clouds that shift stellar motions.
@@
By identifying co-moving structures in a sample of photometrically young stars, SPYGLASS-IV (Kerr et al., 2023) has produced one of the deepest young association surveys to date, revealing 116 young (<50<50 Myr) populations within 1 kpc. Much of this sample is in this largely unknown demographic of low-mass associations, including 10 entirely new populations. These populations are defined by as few as 10 high-confidence photometrically young stars, and as such, they are among the lowest-mass associations ever discovered. They may therefore represent a unique demographic that traces gas structures distinct from the kpc-scale structures that dominate the solar neighborhood (e.g., Lallement et al., 2019; Alves et al., 2020), which could produce unusual initial mass functions (IMFs) and binarity fractions (e.g., Moraux et al., 2007; Conroy & van Dokkum, 2012; Grudic et al., 2023).
@@
In this paper, we perform the first comprehensive study of the ages and demographics of 15 low-mass stellar populations identified in SPYGLASS-IV, revealing new coherent substructures, and providing the high-level overview necessary to guide follow-up dynamical studies. In Section 2, we introduce our Gaia-based dataset alongside new and existing spectroscopic data. We analyze this data in Section 3, assessing stellar membership and identifying substructure. We then compute robust ages for all populations and substructures and measure their masses and demographics in Section 4. We discuss those results in Section 5, before summarizing in Section 6.
@@ 2.1 Initial Population Selection
Our initial membership samples are drawn from SPYGLASS-IV in addition to phase space neighbors, which include early-type stars that are no longer on the pre-main sequence. We select target associations based on a combination of proximity, location, and size, requiring a distance within 350 pc, on-sky position accessible to telescopes at subtropical northern latitudes ($\delta>-30$), and a small population defined by $N\leq 50$ photometrically young stars. We also require that each population has an age solution $\tau<40$ Myr from SPYGLASS-IV to exclude older populations where dynamical ages are less reliable and youth indicators like Lithium are less informative. This produces an initial sample of 19 populations: SCYA-2, SCYA-3, SCYA-26, UPK 88 (SCYA-30), SCYA-35, SCYA-43, Theia 232 (SCYA-54), SCYA-58, Taurus-Orion 1 (SCYA-64), Canis Major North (SCYA-65), Theia 72 (SCYA-66), SCYA-70, SCYA-72, Ophiuchus Southeast (SCYA-75), Aquila East (SCYA-78), SCYA-95, SCYA-97, SCYA-104, and Theia 78 (SCYA-112). We remove SCYA-95, as it shows substantial field contamination that greatly increases the observing time required to spectroscopically survey its members. Finally, we add Theia 98 (SCYA-79), which passes all restrictions except distance, at $d=436$ pc, but has a position and velocity that may suggest dynamical connections to the Cep-Her Complex, alongside Theia 232, making its inclusion useful to assess that possibility in future dynamical studies.
@@ 2.2.2 New Spectroscopy
Gaia ID Assn. SG RA Dec d g M RV (km s-1)aaThe source of the measurement. The codes used are as follows: McDonald Observatory 2.7m Harlan J. Smith Telescope (HJST), Gaia DR3 (GDR3), SDSS APOGEE spectrograph (Wilson et al., 2019) (APOG), SDSS BOSS spectrograph (Smee et al., 2013) (BOSS), DESI DR1 (Collaboration et al., 2025) (DESI), GALAH DR4 (Buder et al., 2024) (GAL4), LAMOST DR9 MRS spectra (Zhang et al., 2021) (LAM9M), LAMOST DR8 LRS parameters of A, F, G, and K stars (LAM8L), LAMOST DR8 Catalog of M Stars (LAM8LMCAT), LAMOST DR8 Value Added Catalog (Wang et al., 2022) (LAM8LV), Famaey et al. (2005) (F5), Torres et al. (2006) (T6), Gontcharov (2006) (G6), Grieves et al. (2018) (G18), Žerjal et al. (2021) (Z21) EWLi (Å)aaThe source of the measurement. The codes used are as follows: McDonald Observatory 2.7m Harlan J. Smith Telescope (HJST), Gaia DR3 (GDR3), SDSS APOGEE spectrograph (Wilson et al., 2019) (APOG), SDSS BOSS spectrograph (Smee et al., 2013) (BOSS), DESI DR1 (Collaboration et al., 2025) (DESI), GALAH DR4 (Buder et al., 2024) (GAL4), LAMOST DR9 MRS spectra (Zhang et al., 2021) (LAM9M), LAMOST DR8 LRS parameters of A, F, G, and K stars (LAM8L), LAMOST DR8 Catalog of M Stars (LAM8LMCAT), LAMOST DR8 Value Added Catalog (Wang et al., 2022) (LAM8LV), Famaey et al. (2005) (F5), Torres et al. (2006) (T6), Gontcharov (2006) (G6), Grieves et al. (2018) (G18), Žerjal et al. (2021) (Z21)bbReported values include the correction for Fe I line deblending described in Section 3.1.5 EW (Å)aaThe source of the measurement. The codes used are as follows: McDonald Observatory 2.7m Harlan J. Smith Telescope (HJST), Gaia DR3 (GDR3), SDSS APOGEE spectrograph (Wilson et al., 2019) (APOG), SDSS BOSS spectrograph (Smee et al., 2013) (BOSS), DESI DR1 (Collaboration et al., 2025) (DESI), GALAH DR4 (Buder et al., 2024) (GAL4), LAMOST DR9 MRS spectra (Zhang et al., 2021) (LAM9M), LAMOST DR8 LRS parameters of A, F, G, and K stars (LAM8L), LAMOST DR8 Catalog of M Stars (LAM8LMCAT), LAMOST DR8 Value Added Catalog (Wang et al., 2022) (LAM8LV), Famaey et al. (2005) (F5), Torres et al. (2006) (T6), Gontcharov (2006) (G6), Grieves et al. (2018) (G18), Žerjal et al. (2021) (Z21) Flags ccFlags include: $V$, velocity membership assessment, where 1 is a kinematic member, 0 is inconclusive, and -1 is a non-member; Li, Lithium membership assessment, where 1 is a member, 0 is inconclusive, and -1 is a non-member; $\alpha, H$$\alpha$ membership assessment, where 0 is inconclusive, and -1 is a non-member; and $F$, General flag, where 1 indicates a star with a resolved companion within 10,000 au in the plane of the sky, 2 indicates a bad broadening function solution, 4 indicates a bimodal line profile likely indicative of spectroscopic binarity, 8 indicates an RUWE>>1.2, indicating likely unresolved binarity, and 16 indicates that the RV recorded was ambiguously attributed to two components of a binary pair. General flags are when multiple are true; for example, flag 6 indicates both flags 2 and 4. Membership ddMembership probabilities provided are: $P_{spat}$, the spatial membership probability, previously referred to as $P_{mem}$ in SPYGLASS-IV, compares number of nearby young and old stars; $P_{sp}$, the spatial-photometric membership probability, defined by correcting the prior on $P_{Age<50Myr}$ from SPYGLASS-IV to align with $P_{spat}$; and $P_{fin}$, the final membership probability, which includes radial velocity and spectroscopic youth indicators
@@ 2.2.3 Data Combination
Following SPYGLASS-V, we use the lowest-uncertainty measurement in cases where multiple sources are available for the same star. We de-prioritize Gaia DR3, RAVE DR6, and LAMOST LRS radial velocities, as these have been shown to have discrepancies from ground based RV measurements of young stars of order 1 km s-1 that are not acknowledged by the reported uncertainties (Kounkel et al., 2023; Kerr et al., 2024). Non-deprioritized observations are always selected over deprioritized sources if both have sub-km s-1 RV measurements. Finally, using the Gaia DR3-2MASS best neighbor cross-match, we add 2MASS K-band photometry to our sample, which will be useful for upcoming Lithium depletion boundary studies (Skrutskie et al., 2006). We compile our complete dataset for all 15 target populations in Table 1.
@@ 3.1 Membership
Our membership assessment in these populations largely follows the probabilistic approach used in SPYGLASS-V, where some steps discussed here are explained in more detail. This approach combines existing spatial and photometric membership assessments with spectroscopic membership indicators like radial velocities and Li EWs to produce aggregate membership probabilities, which we use to separate probable members from background contamination.
@@ 3.1.1 Initial Selection
Many of these populations have a pre-main sequence that separates cleanly from the field. In these cases we use the full SPYGLASS-IV sample without restriction. However, other populations have stellar distributions that are more intertwined with the field, and this is marked by low values of $P_{spatial}$, a metric from SPYGLASS-IV that measures the local fraction of young stars compared to old stars. In these contaminated populations, stars with low $P_{spatial}$ can dominate the color-magnitude diagram (CMD) such that overluminous field stars like binaries fill the gap between the field and pre-main sequence. In these cases, we remove stars with low $P_{spatial}$ to suppress this potentially dominant source of contamination. We therefore apply restrictions to Aries South ($P_{spatial}>0.1$) and Taurus-Orion 1 ($P_{spatial}>0.15$), as well as SCYA-54 and SCYA-79, where we use the $P_{spatial}>0.2$ cut that SPYGLASS-V used in the adjacent and similarly-aged Cep-Her complex. We do not remove stars that were identified as founding photometrically young members of the association by HDBSCAN in SPYGLASS-IV regardless of their $P_{spatial}$. The remaining sample is used for analysis, and comprises 4244 stars across 15 populations.

### v3

@@ Abstract
Recent Gaia-based young stellar association surveys have revealed dozens of low-mass populations that have, until recently, been too small or sparse to detect. These populations represent a largely unstudied demographic with unknown origins, and their relative isolation may minimize gravitational disruptions that impact traceback, making them compelling targets for dynamical studies. In this paper, we survey 15 of these isolated young associations for the first time: Andromeda South (SCYA-97), Aquila East, Aries South (SCYA-104), Cassiopeia East (SCYA-43), Canis Major North, Leo Central (SCYA-2), Leo East (SCYA-3), Theia 72, Ophiuchus Southeast, Scutum North (SCYA-70), Taurus-Orion 1 (TOR1), Theia 78, Vulpecula East (UPK 88), SCYA-54, and SCYA-79. By combining Gaia astrometry and photometry with new ground-based spectroscopic measurements, we assess the membership of each population, search for substructure, analyze their demographics, and compute ages. We find that the smallest populations in our sample contain <20<20 $M_{\odot}$ of stellar mass, making them the smallest associations ever detected. Four host substantial substructure, including TOR1, where we discover TOR1B, a new $M_{\odot}$ association with radial velocities inconsistent with an origin in the parent complex. Using PARSEC isochrones, we produce self-consistent ages for all populations supported by dynamical and lithium depletion ages, which range from 6.9 $\pm$ 0.5 Myr in to 42.8 $\pm$ 2.4 in AndS. Our results provide the first detailed overview of the properties of these populations, characterizing a largely unknown category of young associations that may have an important role in tracing the processes that guide local star formation.
@@
facilities: Gaia, McDonald Observatory 2.7m Harlan J. Smith Telescope, Sloan Digital Sky Survey BOSS and APOGEE spectrographs at APO and LCOsoftware: astropy (Astropy13; Astropy18; astropy22), numpy (numpy), pandas (pandas), LMFIT (lmfit) saphires (Tofflemire19), matplotlib (Hunter07)
@@ 1 Introduction
Most nearby young stars are not found in isolation, but rather in co-moving star clusters and associations (deZeeuw99; Krumholz19). These populations inherit motions from their parent molecular cloud, and as such their positions, velocities and ages produce a record of the positions and dynamics of star-forming clouds long after their dispersal (e.g. Galli23; Swiggum24). On small scales, young stellar populations are used to detect star formation patterns, such as sequential and triggered star formation, where feedback from early-forming stars or gas collisions may compress gas and trigger new stellar generations (Tan00; Inutsuka15; Posch23). On larger scales, the locations of these clouds at formation collectively trace the sites of dense gas in our Galaxy over time. An improved census of recent star formation will therefore reveal not just local star formation patterns, but also kpc-scale gas patterns that trace larger structures like spiral arms.
@@
The data releases of the Gaia spacecraft are making the local star formation record more complete than ever before (GaiaMission; GaiaDR322). Recent Gaia-based surveys have substantially deepened our record of nearby associations, ranging from large, heavily substructured stellar complexes to small, seemingly isolated associations (e.g. Kounkel19; Hunt23). Many of the larger populations in the solar neighborhood have already been studied, often revealing extensive substructure and multi-generational star formation patterns (e.g., CantatGaudin19; Kerr21; Briceno23; Ratzenbock23; SanchezSanjuan24; Kerr24).
@@
However, associations containing less than a few hundred solar masses often lack dedicated studies of their structure, demographics, and ages, especially outside of the nearest 100 pc (deZeeuw99; Gagne18BXIII). As such, we know little about their typical dynamics and star formation patterns. We also know little about whether they emerge from the kpc-scale gas structures in the solar neighborhood or any clouds that may have preceded their assembly, or whether they form in truly isolated environments. If their formation is isolated, they may be better-suited for precise dynamical studies, having had minimal interactions with particularly massive nearby gas clouds that shift stellar motions.
@@
By identifying co-moving structures in a sample of photometrically young stars, SPYGLASS-IV (Kerr23) has produced one of the deepest young association surveys to date, revealing 116 young (<50<50 Myr) populations within 1 kpc. Much of this sample is in this largely unknown demographic of low-mass associations, including 10 entirely new populations. These populations are defined by as few as 10 high-confidence photometrically young stars, and as such, they are among the lowest-mass associations ever discovered. They may therefore represent a unique demographic that traces gas structures distinct from the kpc-scale structures that dominate the solar neighborhood (e.g., Lallement19; Alves20), which could produce unusual initial mass functions (IMFs) and binarity fractions (e.g., Moraux07; Conroy12; Grudic23).
@@
In this paper, we perform the first comprehensive study of the ages and demographics of 15 low-mass stellar populations identified in Kerr23, revealing new coherent substructures, and providing the high-level overview necessary to guide follow-up dynamical studies. In Section 2, we introduce our Gaia-based dataset alongside new and existing spectroscopic data. We analyze this data in Section 3, assessing stellar membership and identifying substructure. We then compute robust ages for all populations and substructures and measure their masses and demographics in Section 4. We discuss those results in Section 5, before summarizing in Section 6.
@@ 2.1 Initial Population Selection
Our initial membership samples are drawn from Kerr23, which identifies photometrically young stars using a Bayesian framework, and then applies the HDBSCAN clustering algorithm (McInnes2017) to detect overdensities of ten or more photometrically young Gaia stars that mark the site of an association. The resulting catalog provides both the photometrically young stars with robust membership used to define each group, in addition to phase space neighbors, which include early-type stars that are no longer on the pre-main sequence. We select target associations based on a combination of proximity, location, and size, requiring a distance within 350 pc, on-sky position accessible to telescopes at subtropical northern latitudes ($\delta>-30$), and a small population defined by $N\leq 50$ photometrically young stars. We also require that each population has an age solution $\tau<40$ Myr from Kerr23 to exclude older populations where dynamical ages are less reliable and youth indicators like Lithium are less informative. This produces an initial sample of 19 populations: SCYA-2, SCYA-3, SCYA-26, UPK 88 (SCYA-30), SCYA-35, SCYA-43, Theia 232 (SCYA-54), SCYA-58, Taurus-Orion 1 (SCYA-64), Canis Major North (SCYA-65), Theia 72 (SCYA-66), SCYA-70, SCYA-72, Ophiuchus Southeast (SCYA-75), Aquila East (SCYA-78), SCYA-95, SCYA-97, SCYA-104, and Theia 78 (SCYA-112). We remove SCYA-95, as it shows substantial field contamination that greatly increases the observing time required to spectroscopically survey its members. Finally, we add Theia 98 (SCYA-79), which passes all restrictions except distance, at $d=436$ pc, but has a position and velocity that may suggest dynamical connections to the Cep-Her Complex, alongside Theia 232, making its inclusion useful to assess that possibility in future dynamical studies.
@@ 2.2.2 New Spectroscopy
Gaia ID Assn. SG RA Dec d g M RV (km s-1)aaThe source of the measurement. The codes used are as follows: McDonald Observatory 2.7m Harlan J. Smith Telescope (HJST), Gaia DR3 (GDR3), SDSS APOGEE spectrograph (APOGEE) (APOG), SDSS BOSS spectrograph (BOSS) (BOSS), DESI DR1 (DESIDR1) (DESI), GALAH DR4 (Buder24) (GAL4), LAMOST DR9 MRS spectra (LAMOSTSCRV) (LAM9M), LAMOST DR8 LRS parameters of A, F, G, and K stars (LAM8L), LAMOST DR8 Catalog of M Stars (LAM8LMCAT), LAMOST DR8 Value Added Catalog (LAMOSTDR8VAC) (LAM8LV), Famaey05 (F5), Torres06 (T6), Gontcharov06 (G6), Grieves18 (G18), Zerjal21 (Z21) EWLi (Å)aaThe source of the measurement. The codes used are as follows: McDonald Observatory 2.7m Harlan J. Smith Telescope (HJST), Gaia DR3 (GDR3), SDSS APOGEE spectrograph (APOGEE) (APOG), SDSS BOSS spectrograph (BOSS) (BOSS), DESI DR1 (DESIDR1) (DESI), GALAH DR4 (Buder24) (GAL4), LAMOST DR9 MRS spectra (LAMOSTSCRV) (LAM9M), LAMOST DR8 LRS parameters of A, F, G, and K stars (LAM8L), LAMOST DR8 Catalog of M Stars (LAM8LMCAT), LAMOST DR8 Value Added Catalog (LAMOSTDR8VAC) (LAM8LV), Famaey05 (F5), Torres06 (T6), Gontcharov06 (G6), Grieves18 (G18), Zerjal21 (Z21)bbReported values include the correction for Fe I line deblending described in Section 3.1.5 EW (Å)aaThe source of the measurement. The codes used are as follows: McDonald Observatory 2.7m Harlan J. Smith Telescope (HJST), Gaia DR3 (GDR3), SDSS APOGEE spectrograph (APOGEE) (APOG), SDSS BOSS spectrograph (BOSS) (BOSS), DESI DR1 (DESIDR1) (DESI), GALAH DR4 (Buder24) (GAL4), LAMOST DR9 MRS spectra (LAMOSTSCRV) (LAM9M), LAMOST DR8 LRS parameters of A, F, G, and K stars (LAM8L), LAMOST DR8 Catalog of M Stars (LAM8LMCAT), LAMOST DR8 Value Added Catalog (LAMOSTDR8VAC) (LAM8LV), Famaey05 (F5), Torres06 (T6), Gontcharov06 (G6), Grieves18 (G18), Zerjal21 (Z21) Flags ccFlags include: $V$, velocity membership assessment, where 1 is a kinematic member, 0 is inconclusive, and -1 is a non-member; Li, Lithium membership assessment, where 1 is a member, 0 is inconclusive, and -1 is a non-member; $\alpha, H$$\alpha$ membership assessment, where 0 is inconclusive, and -1 is a non-member; and $F$, General flag, where 1 indicates a star with a resolved companion within 10,000 au in the plane of the sky, 2 indicates a bad broadening function solution, 4 indicates a bimodal line profile likely indicative of spectroscopic binarity, 8 indicates an RUWE>>1.2, indicating likely unresolved binarity, and 16 indicates that the RV recorded was ambiguously attributed to two components of a binary pair. General flags are when multiple are true; for example, flag 6 indicates both flags 2 and 4. Membership ddMembership probabilities provided are: $P_{spat}$, the spatial membership probability, previously referred to as $P_{mem}$ in Kerr23, compares number of nearby young and old stars; $P_{sp}$, the spatial-photometric membership probability, defined by correcting the prior on $P_{Age<50Myr}$ from Kerr23 to align with $P_{spat}$; and $P_{fin}$, the final membership probability, which includes radial velocity and spectroscopic youth indicators
@@ 2.2.3 Data Combination
Following Kerr24, we use the lowest-uncertainty measurement in cases where multiple sources are available for the same star. We de-prioritize Gaia DR3, RAVE DR6, and LAMOST LRS radial velocities, as these have been shown to have discrepancies from ground based RV measurements of young stars of order 1 km s-1 that are not acknowledged by the reported uncertainties (Kounkel23; Kerr24). Non-deprioritized observations are always selected over deprioritized sources if both have sub-km s-1 RV measurements. Finally, using the Gaia DR3-2MASS best neighbor cross-match, we add 2MASS K-band photometry to our sample, which will be useful for upcoming Lithium depletion boundary studies (Skrutskie06). We compile our complete dataset for all 15 target populations in Table 1.
@@ 3.1 Membership
Our membership assessment in these populations largely follows the probabilistic approach used in Kerr24, where some steps discussed here are explained in more detail. This approach combines existing spatial and photometric membership assessments with spectroscopic membership indicators like radial velocities and Li EWs to produce aggregate membership probabilities, which we use to separate probable members from background contamination.
@@ 3.1.1 Initial Selection
Many of these populations have a pre-main sequence that separates cleanly from the field. In these cases we use the full Kerr23 sample without restriction. However, other populations have stellar distributions that are more intertwined with the field, and this is marked by low values of $P_{spatial}$, a metric from Kerr23 that measures the local fraction of young stars compared to old stars. In these contaminated populations, stars with low $P_{spatial}$ can dominate the color-magnitude diagram (CMD) such that overluminous field stars like binaries fill the gap between the field and pre-main sequence. In these cases, we remove stars with low $P_{spatial}$ to suppress this potentially dominant source of contamination. We therefore apply restrictions to Aries South ($P_{spatial}>0.1$) and Taurus-Orion 1 ($P_{spatial}>0.15$), as well as SCYA-54 and SCYA-79, where we use the $P_{spatial}>0.2$ cut that Kerr24 used in the adjacent and similarly-aged Cep-Her complex. We do not remove stars that were identified as founding photometrically young members of the association by HDBSCAN in Kerr23 regardless of their $P_{spatial}$. The remaining sample is used for analysis, and comprises 4244 stars across 15 populations.